Im only searching for this question "which of the following statements is true" I have done the rest of the problem.

Which of the following statement(s) is/are true?

Revenues decreased Rudy Gandolfi’s equity, and expenses and dividends increased it.checkbox unchecked1 of 5
Interest is a "cost" of doing business.checkbox unchecked2 of 5
Dividend is a "cost" of doing business.checkbox unchecked3 of 5
Cash balance will be affected while paying off an accounts payable and collecting an accounts receivable.checkbox unchecked4 of 5
When money is borrowed from the bank, the total assets and the total liabilities will increase.

Rudy Gandolfi owns and operates Rudy’s Furniture Emporium Inc. The balance sheet totals for assets, liabilities, and stockholders’ equity at August 1, 2019, are as indicated. Described here are several transactions entered into by the company throughout the month of August. Required: Indicate the amount and effect of each transaction on total assets, total liabilities, and total stockholders' equity, and then compute the new totals for each category. The first transaction is provided as an illustration. Indicate the financial statement effect. How much were total revenues and total expenses during August? What was the amount of net income (or loss) during August? What were the net changes during the month of August in total assets, total liabilities, and total stockholders' equity? Which of the following statement(s) is/are true?

a. Assets = Liabilities + Stockholder's equity
August 1,2019 total 700000 500000 200000
August 3, borrowed $ 50000 in cash from the bank 50000 50000 0
New totals 750000 550000 200000
August 7,brought merchandise inventory valued at $ 75000 on account 75000 75000 0
New totals 825000 625000 200000
August 10,paid $25000 cash for operating expenses -25000 0 -25000
New totals 800000 625000 175000
August 14,received $120000 in cash from sales of merchandise that had cost $ 72000
Effect of sales revenue 120000 0 120000
Effect of cost of goods sold -72000 0 -72000
New totals 848000 625000 223000
August 17,paid $ 60000 owed on accounts payable -60000 -60000 0
New totals 788000 565000 223000
August 21,collected $ 44000 of accounts receivable 44000 0 0
-44000 0 0
New totals 788000 565000 223000
August 24,repaid $ 30000 to the bank plus $ 1000 interest -31000 -30000 -1000
New totals 757000 535000 222000
August 29,Paid Rudy Gandolfi a cash dividend of $ 15000 -15000 0 -15000
August 31,2019 totals 742000 535000 207000
b. Total revenue:
Date Description $
Aug 14. Sale of merchandise 120000
Total 120000
Total expenses:
Date Description $
Aug 10. Operating expenses 25000
Aug 14. Cost of goods sold 72000
Aug 24. Interest expenses 1000
Total 98000
Net income=Total revenue-Total expenses=120000-98000=$ 22000
c. Net changes=August 31,2019 totals-August 1,2019 totals
Total assets,
Net changes=742000-700000=$ 42000
Total liabilities,
Net changes=535000-500000=$ 35000
Total stockholder's equity,
Net changes=207000-200000=$ 7000
Answer #1

Sub Part 1

Date Particulars Assets = Liability + Equity Reasons
1-Aug-2019 Opening Balance 700000 = 500000 + 200000
3-Aug-2019 Borrowings from Bank 50000 = 50000 + 0 Money Borrowed from bank
1. Cash being a asset Increases when received
2. Borrowings-->> Liabilities Increases
7-Aug-2019 Inventory Purchased on account 75000 = 75000 + 0 Inventory purchasd for Credit/Loan
1. Inventory being a asset Increases
2. Loan being a liability Increases.
10-Aug-2019 cash for operating expenses -25000 = 0 + -25000 Cash being an asset decreases when paid and Expenses increases.
14-Aug-2019 Sale of merchandise 120000 = 0 + 120000 1. Cash being an asset When received from sale increases
2. Inventories being an asset decreases when sold, and being carried at cost they are to be reduced to such Extent only
3. Profit of 120000-72000 to be considered in equity
Effect of cost of goods sold -72000 = 0 + -72000
17-Aug-2019 owed on accounts payable -60000 = 0 + -60000 1. Repayment of Liability
Reduces Borrowings and cash Balance
21-Aug-2019 Collected from Account receivable 44000 = 0 + 0 Amount collected from account receivable
1. Cash Being an asset increases when received
2. Account receivable being a asset decreases.
21-Aug-2019 Collected from Account receivable -44000 = 0 + 0
24-Aug-2019 August 24,repaid $ 30000 to the bank plus $ 1000 interest -31000 = -30000 + -1000 1. Cash Cash being an asset When paid decreases.
2. Borriowing being liability reduces when paid. And shall be reduced to the extent of prinicipal amount. Balance is reduced from Equity.Since interest is in the nature of Expense
29-08-2019 August 29,Paid Rudy Gandolfi a cash dividend of $ 15000 -15000 = + -15000 1. Cash Cash being an asset When paid decreases.
2. Dividend is in the nature of expense, to reduced from equity.
TOTAL 742000 = 595000 + 147000

Sub Part 2

S.No Partculars Debit Credit
1 Sale of merchandise 120000
2 Cost of Goods sold -72000
3 Operating Expenses -25000
4 Bank Interest -1000 -98000
5 Net Profit 22000
6 Less: Dividend -15000
5 Profit after Dividend 7000

Sub Part 3

S.No Partilculars Asset = Liabillity + Equity
1 Closing Balance as per Sub part 1 742000 = 595000 + 147000
2 Opening Balance as per question 700000 = 500000 + 200000
3 Net Changes 42000 = 95000 + -53000

Notes:
1. This statement is true, Revenues decreased Rudy Gandolfi’s equity, and expenses and dividends increased it refer sub part3

2.Interest is a cost for doing a business - Yes this statement is true

3. Dividend is a "cost" of doing business - Dividend is reward to shareholders, It is a cost to company - This statement is true.

4. Cash balance will be affected while paying off an accounts payable and collecting an accounts receivable - Yes this statement is true Explanation: Cash Being an asset increases when received and decreases when paid.

5.When money is borrowed from the bank, the total assets and the total liabilities will increase. - This statement is true Explanation: When Cash Recived it increases (Asset side) and Loan increases (Liability side)